(b) 10 points Find all complex numbers z satisfying 28 – 324 – 4 = 0.
Answer #1

Z=-2 z8-324-4=0. or. Z®-42+2-420. av. Z[z-4) +(z-4)=0. ew. (z-4) (zº+1) =0. wr. (772)(z?-2) (z4+1) = 0. jis 7+2=0. | 77